OpenStudy (anonymous):

what is answer for 16 x 10^0 (10 to the power of 0)

jimthompson5910 (jim_thompson5910):

10^0 = 1 So 16 x 10^0 = 16 x 1 = 16

OpenStudy (anonymous):

10^0 = 1 so 16 X 1 = 16